# User Simulation for Evaluating Information Access Systems

### v1

@@ 6.9 Summary and Future Challenges
Conversational AI broadly refers to systems that are capable of natural language understanding and responding in a way that mimics human dialogue, thereby providing an interactive and engaging experience for users. Traditionally, conversational systems are categorized as being goal-driven (or task-oriented), aiming to assist users to complete some specific task, or non-goal-driven (also known as chatbots), aiming to carry on an extended conversation (“chit-chat”), usually with the purpose on entertainment [Chen:2017:SIGKDD, Jurafsky:2023:book]. Our interest in this book lies in the former category, that is, conversational systems that have a clear goal: assisting the user in completing some task with an underlying information need, which can be satisfied through a conversation. This includes the tasks of search, recommendation, and question answering (QA), which are collectively referred to as the problem of conversational information access (CIA) [Balog:2021:DESIRES] (a.k.a. conversational information seeking [Zamani:2023:arXiv]), since the boundaries between the different tasks are often blurred. Conversational systems, unlike traditional search engines, are also naturally suited to address exploratory search needs [Zamani:2023:arXiv].
Even though interactions between a user and a traditional search and recommender system are a form of conversation, the users of those systems have access to only a limited set of actions allowed by a system’s user interface (clicking, scanning, rating, etc.). In contrast, in a truly conversational setting, the user can freely express their request and feedback in natural language. This extends beyond identifying slot-value pairs, typical of early dialogue systems (e.g., departure and destination city/airport, departure and return dates, etc. for travel booking). Since user intents are no longer captured as explicit actions, such as querying or clicks, they need to be inferred from free text, which requires more advanced natural language understanding capabilities. Even in seemingly straightforward cases, such as yes/no questions, humans might answer indirectly. For example, when asked about their interest in an evening activity, the user might respond with “I’d prefer to go to bed” instead of simply saying “no” [Louis:2020:EMNLP].
Another crucial factor that sets CIA apart from traditional search and recommender systems is the notion of initiative. The interaction in traditional search systems is driven by the user, i.e., the user issues a request (query) and the system responds with a ranked list of results, also known as the query-response paradigm. In recommender systems, on the other hand, the system takes initiative by offering suggestions to the user based on their history. Mixed-initiative refers to a conversational paradigm where the user and system both actively participate in addressing the user’s information need. Mixed-initiative has been identified as a critical component of conversational information access systems [Radlinski:2017:CHIIR, Trippas:2018:CHIIR, Wadhwa:2021:DESIRES]. It enables systems to engage with users in natural language, for example, to ask clarification questions, solicit feedback, or to learn about the user’s preferences in specific contexts. The mixed-initiative paradigm in CIA clearly presents special challenges for user simulation.
Overall, conversational systems can reduce effort for the user from a cognitive perspective, but there is also more potential for frustration [Luger:2016:CHI], as the user has less direct control over the behaviour of the system, making the problem of user simulation in a conversational setting less well defined as compared with a search engine or recommender system.
What makes user simulation for conversational systems particularly challenging is that simulators need to have similarly advanced natural language understanding and generation capabilities as CIA systems. Results are no longer restricted to (a ranked list of) items from a predefined set (a document collection or a catalog of items), but could be text of arbitrary length (possibly containing semi-structured elements, such as lists, tables, or images embedded in case of a chat-based user interface), a question posed to the user, or the system expressing in some way that it could not understand or help with the request. While query generation can build on work done in the context of search engines (cf. Section 6.2), there, queries are typically sequences of keywords and not full-fledged natural language questions. Further, other actions, such as clicks or ratings, are no longer captured as discrete events, but need to be articulated as nuanced natural language expressions.
Because of significant differences between conversational systems and traditional search/recommender systems, this chapter is organized differently from the previous one. In Section 7.1 we conceptualize conversational information access in terms of intents and dialogue structure. Next, we discuss two fundamentally different simulator architectures in Section 7.2, which correspond to intent-based and text-based generation of user utterances. There is a solid body of work within dialogue systems research on simulating user decisions, which we cover in Section 7.3. This is followed by the discussion of simulation approaches developed specifically for conversational information access in Section 7.4. Section 7.5 reviews toolkits and resources, followed by a discussion on how simulators themselves can be evaluated in Section 7.6. It is worth emphasizing that user simulation for CIA evaluation is very much an area of ongoing work, where notable advancements are being made in terms of conversational models as well as evaluation measures and methodologies. We identify and discuss some of the future challenges in Section 7.7.
7.1 Conceptualization of Conversational Information Access
@@
We assume that a conversation happens between a user and a system, who take turns. Following existing research, we assume the modality to be text, either written or spoken (and transcribed), noting that other modalities are also receiving increased attention. A dialogue, thus, is a sequence of turns, where each turn is a natural language utterance from either the user or the system. In terms of functionality, we expect conversational information access systems to (1) support multiple user goals, including search, recommendation and exploratory information gathering, (2) learn user preferences and personalize responses accordingly, and (3) be capable of taking initiative [Balog:2021:DESIRES].
It is important to note that while the goals that CIA systems should support are clear, much of the envisioned functionality, as of today, mostly exists only on a conceptual level. Radlinski:2017:CHIIR identify desirable properties of conversational search system, which include the following: (1) user revealment (enabling the user to express/discover their information need), (2) system revealment (enabling the system to reveal its capabilities and corpus), (3) mixed initiative (allowing both the user and system to take initiative), (4) memory (the ability to reference and incorporate past statements), and (5) set retrieval (considering utility over sets of complementary items). Trippas:2018:CHIIR provide a different high-level formalization, distinguishing three stages of the conversational search process: (1) query formulation, (2) search results exploration, and (3) query re-formulation. In order to simulate how users interact with CIA systems, we need to abstract out both low-level actions and decision points as well as higher-level dialogue patterns that manifest during conversational interactions.
7.1.1 Dialogue Acts
@@
The notion of dialogue act originates from (spoken) dialogue systems research and is meant to represent the function or high-level intention of an utterance in a dialogue, for example, whether it is a question, offer, suggestion, etc. [McTear:2016:book]. Dialogue acts are typically represented as tuples, consisting of an intent and, optionally, a number of slot-value pairs (e.g., AFFIRM or INFORM(a=x,b=y,...)). [^1] We note that concept of dialogue act is not used consistently the literature; in some cases it refers to the intent, while in other cases it means the intent along with the associated slot-value pairs. For clarity, we use intent in the former and dialogue act in the latter case. When the distinction is not important or the sense is clear from the context, we may simply write action. The set of dialogue acts needs to be designed specific to the objectives of the dialogue application. Over the years, various taxonomies have been developed. For example, the Dialogue State Tracking Challenge [Williams:2016:DD] has defined user and system actions for task-oriented dialogue systems (bus, restaurant, and tourist information). The international standard ISO 24617-2 defines communication functions for general dialogue and also provides a formal language, the Dialogue Act Markup Language (DiAML), for expressing dialogue annotations [Bunt:2017:chapter, Bunt:2020:LREC].

### v2

@@ 6.7 Summary and Future Challenges
Conversational AI broadly refers to systems that are capable of natural language understanding and responding in a way that mimics human dialogue, thereby providing an interactive and engaging experience for users. Traditionally, these systems are categorized as being goal-driven (or task-oriented), aiming to assist users to complete some specific task, or non-goal-driven (also known as chatbots), aiming to carry on an extended conversation (“chit-chat”), usually with the purpose on entertainment [Chen:2017:SIGKDD, Jurafsky:2023:book]. Our interest in this book lies in the former category, that is, conversational systems that have a clear goal: assisting the user in completing some task with an underlying information need, which can be satisfied through natural language dialogue. This includes the tasks of search, recommendation, and question answering, collectively referred to as the problem of conversational information access (CIA) [Balog:2021:DESIRES]. CIA can be seen as a specific form of task-oriented dialogue (TOD), where the user’s goal can, for instance, be to satisfy some ad-hoc information need, find an item matching their preferences, or learn about a topic. What makes user simulation for conversational systems particularly challenging is that a simulator also needs to possess advanced natural language understanding and generation capabilities (cf. Section 2.1.3). Consequently, constructing a fully functional user simulator is akin to building an independent conversational agent.
We begin by establishing basic concepts in Section 7.1. Simulator architectures, which closely mirror the architectural principles of TOD systems, are introduced in Section 7.2. User simulation for TOD has a longstanding history in dialogue systems research and has recently gained increased interest within the natural language processing community, with a growing number of approaches utilizing pre-trained large language models (LLMs). We review simulation for TOD systems in Section 7.3. Although conversational information access can be seen as a form of task-oriented dialogue, it goes beyond the identification and filling of slot-value pairs, typical of traditional TOD systems (e.g., departure and destination city/airport, departure and return dates, etc. for travel booking). Therefore, in Section 7.4, we discuss how CIA can be conceptualized in terms of intents and dialogue structure. This is followed by the discussion of simulation approaches developed specifically for conversational search and conversational recommendation in Sections 7.5 and 7.6, respectively. It is worth emphasizing that user simulation for CIA evaluation is very much an area of ongoing work, where notable advancements are being made in terms of conversational models as well as evaluation measures and methodologies. We identify and discuss some of the future challenges in Section 7.7.
7.1 Preliminaries
@@
We assume that a conversation happens between a user and a system, who take turns. Following existing research, we assume the modality to be text, either written or spoken (and transcribed), noting that other modalities are also receiving increased attention. A dialogue, thus, is a sequence of turns, where each turn $t$ is a natural language utterance from either the user $x_{t}^{u}$ or the system $x_{t}^{s}$: $x_{0}^{s}\rightarrow x_{0}^{u}\rightarrow x_{1}^{s}\rightarrow x_{1}^{u}\rightarrow\dots\rightarrow x_{t}^{s}\rightarrow x_{t}^{u}$.
@@
The notion of dialogue act originates from (spoken) dialogue systems research and is meant to represent the function or high-level intention of an utterance in a dialogue, for example, whether it is a question, offer, suggestion, etc. [McTear:2016:book]. Dialogue acts are typically represented as tuples, consisting of an intent and, optionally, a number of slot-value pairs (e.g., AFFIRM or INFORM(a=x,b=y,...)). [^1] We note that concept of dialogue act is not used consistently in the literature; in some cases it refers to the intent, while in other cases it means the intent along with the associated slot-value pairs. For clarity, we use intent in the former and dialogue act in the latter case. When the distinction is not important or the sense is clear from the context, we may simply write action. Formally, a dialogue act is a tuple $a=(I,V)$, where $I\in\mathcal{I}$ is an intent from the set of possible dialogue intents and $V$ is a list of slot-value pairs $V=\langle(slot_{1},value_{1}),\dots,(slot_{n},value_{n})\rangle$. Human-machine dialogue can then be represented semantically as a sequence dialogue acts by the system ($a_{i}^{s}$) and the user ($a_{i}^{u}$) as they take turns: $a_{0}^{s}\rightarrow a_{0}^{u}\rightarrow a_{1}^{s}\rightarrow a_{1}^{u}\rightarrow\dots\rightarrow a_{t}^{s}\rightarrow a_{t}^{u}$.
The space of possible intents and slots need to be defined specific to the objectives of the dialogue application. Over the years, various taxonomies have been developed. For example, the Dialogue State Tracking Challenge [Williams:2016:DD] has defined user and system actions for task-oriented dialogue systems (bus, restaurant, and tourist information). The international standard ISO 24617-2 defines communication functions for general dialogue and also provides a formal language, the Dialogue Act Markup Language (DiAML), for expressing dialogue annotations [Bunt:2017:chapter, Bunt:2020:LREC].
7.2 Simulator Architectures
The main challenge in simulating users in a conversational setting is that the simulator needs to have advanced natural language understanding and generation capabilities, much like that of a conversational system. Hence, adopting the architecture of task-oriented dialogue systems is a natural approach in the construction of user simulators. Accordingly, two broad categories of simulator architectures can be distinguished: modular (or pipelined) and end-to-end systems. The difference between these two is more fundamental than just an architectural choice, it is also a question of granularity: modular simulators model user responses semantically on the level of dialogue acts, and then generate the corresponding natural language utterances. End-to-end systems, on the other hand, operate on the utterance level, i.e., generate textual responses directly. While end-to-end systems might yield more fluent dialogues, they do not allow us to interpret user behaviour, and therefore are not that well suited for our purposes of evaluation. Parameterization of such systems with meaningful user variables would be desirable in general, but this requires modeling a user beyond their surface-level utterances in order to capture latent user states using more interpretable models.
7.2.1 Modular Systems
Dialogue management is responsible for maintaining the dialogue state and determining the next user action based on the system’s utterance and the goals, knowledge, and preferences of the individual. The dialogue state in traditional task-oriented dialogue systems is based around the notion of a semantic frame, which is a knowledge structure consisting of a collection of slots that together specify what the system needs to know to complete a given task [Jurafsky:2023:book]. [^2] In spoken dialogue systems, a considerable research effort is centered on dealing with the uncertainty resulting from noise in automatic speech recognition. There, it is common for the dialogue system to maintain multiple hypotheses of the true dialog state [Williams:2016:DD]. However, for text-based interactions, this is typically not a concern. For user simulation, the dialogue state needs to capture dialogue context from the user’s perspective. This includes representing the current information need, the dialogue history, and the system’s last action. The dialogue state may be represented explicitly, in a directly observable and interpretable way, such as previous dialogue acts and the status of various slots, or as a latent representation in an embedding space. The dialogue policy determines how the user should respond, i.e., the user intent and associated parameters; this might be influenced by the individual characteristics of the user. To ensure that the simulator maintains a dialogue strategy, a commonly used approach is an agenda-based simulator; it maintains an agenda or a stack of actions and determines which action to execute at each turn. Alternatively, the policy may be learned directly from data. We discuss these alternatives for TOD in more detail in Section 7.3.
Natural language generation (NLG) is concerned with turning the generated response from a structured representation (dialogue act) into natural language. At its simplest form, this is based on templates [Gur:2018:SLT, Shi:2019:EMNLP]. Other alternatives include retrieval-based methods [Wu:2017:ACL, Hu:2014:NIPS], neural text generation models [Wen:2015:EMNLP, Tseng:2021:ACL], and hybrid solutions [Li:2017:arXiv, Sun:2023:TOIS]. Natural language generation can also consult the user modeling component to produce outputs that are customized based on individual characteristics.
User modeling focuses on capturing the characteristics of individuals that would influence how they interact with the system. These could include, among others, information about the user’s goal, knowledge, preferences, personal characteristics (e.g., patience), and beliefs about the system (e.g., based on past interactions). In case of information access tasks, one can often identify commonalities (similar user state variables) between user modeling in a conversational system and traditional search/recommender systems. However, both the information available for inferring a user’s state and the observable actions are different. In particular, for conversational systems, we would need to model how a user’s latent state influences the natural language utterances generated by the user as well as infer a user’s latent state based on the natural language system responses.
7.2.2 End-to-End Systems
Figure 7.2: End-to-end user simulator.
More recently, there have been numerous attempts to build end-to-end trainable task-oriented dialogue systems, leveraging advances in end-to-end neural generative models [Gao:2019:FnTIR, Ni:2022:AIR]. The main benefits over modular systems include the possibility of global optimization and easier adaptation to new domains [Chen:2017:SIGKDD]. End-to-end architectures may also be used for constructing user simulators, by casting the generation of user responses as a sequence-to-sequence (or source-to-target transduction) problem: given a dialogue history up to the given turn as input, generate the user’s natural language response as the output sequence. [^3] As noted above for modular architectures, in end-to-end simulation the NLU component may also be omitted [Lin:2022:SIGDIAL]. Instead, the dialogue act corresponding to the system utterance is taken as input, effectively equipping the simulated user with perfect language understanding capabilities. This problem can be effectively addressed by transformer-based natural language generation models [Tseng:2021:ACL, Lin:2022:SIGDIAL]. However, as indicated earlier, such model-based end-to-end approaches are more suitable for other uses of simulation, such as RL-based learning [Lin:2022:SIGDIAL, Tseng:2021:ACL] or data augmentation [Wan:2022:EMNLP, Mohapatra:2021:EMNLP, Kim:2021:ACL]. For simulation-based evaluation, a more controllable generation of user responses is required, which (currently) cannot be guaranteed with end-to-end systems.
7.3 User Simulation for Task-Oriented Dialogue
In this section, we discuss work on simulating users in the context of traditional (“slot-filling”) task-oriented dialogue systems. Such systems are built for specific domains, each characterized by a domain ontology that describes the specific intents, slots, and entities that can be talked about. The supported conversational functionality is limited to the user specifying their constraints in terms of informable slots ($C$) and requesting information on requestable slots ($R$) to find matching entities in a knowledge base. This is analogous to some scenarios considered within conversational information access, such as finding items where the user already has a reasonably concrete idea of attributes and the constraints on them, or looking up known items or their properties. However, CIA systems ultimately support a much broader array of information needs. This is discussed further in Section 7.4.
Below, we present a selection of influential approaches, categorized into model-based and data-driven approaches, with respect to the modeling of user decisions. Within each category, the approaches are discussed in a roughly chronological order, highlighting the progression of methods and the ways in which subsequent work builds upon and is influenced by earlier research.
7.3.1 Model-Based Approaches
Model-based approaches are characteristic of modular simulator architectures, where the user’s response is governed by a dialogue policy (a.k.a. dialogue strategy). [^4] It is not to be confused by the state model and policy of the conversational agent; while task-oriented dialogue systems are often built on top MDPs, those are different for the system and the (simulated) user. The policy $\pi$ determines what action $a^{u}_{t+1}$ the user should take next, given the dialogue history up to turn $t$. A main challenge here is that the simulator should give responses that are coherent across multiple conversational turns and are driven by some user goal. This can naturally be modeled in the MDP framework. Statistical user modeling techniques “exploit the Markov assumption that the next user action can be predicted based on some representation of the current state of the dialogue” [Schatzmann:2006:KER]. As the user and system take turns, the simulator receives the corresponding dialogue acts ($a^{u}_{t}$ for user and $a^{s}_{t}$ for system), and updates the state:
N-grams Models
$\pi(s_{t})=P(a^{u}_{t+1}|a^{s}_{t})~{},$ (7.3)
where the conditional probabilities are estimated based on an annotated corpus of human-machine dialogues. This model does not have any information about the user’s goal and does not place any constraints on the simulated user behaviour, therefore fails to produce realistic dialogues. Subsequent work tried to correct for it by estimating a selected set of simpler probabilities corresponding to “sensible” pairs of system and user actions (specifically, responding to greetings, constraining questions, and relaxation prompts), while other are assumed to be zero, thereby setting some constraints on the flow of the dialogue [Levin:2000:TSAP]. [^5] In principle, all the selected probabilities could be estimated from a dialogue corpus, but in practice, the probabilities corresponding to constraining or relaxing questions are set manually in [Levin:2000:TSAP] as they were not present in the corpus used. While this yields somewhat more realistic dialogues, the consistency between user responses across the dialogue is still not guaranteed. Also, simulated dialogues might run infinitely long if the user keeps changing their answer to the same question.
Goal-directed User Model with Memory
Agenda-based Simulator
In [Schatzmann:2007:NAACL], the agenda is initialized with all goal constraints set to INFORM acts and all goal requests set to REQUEST acts; the BYE act is added at the bottom of the agenda to close the dialogue. For the above example, this corresponds to:
$P(s_{t+1}|s_{t},a_{t}^{u})$ $=P(A_{t+1},G_{t+1}|A_{t},G_{t},a_{t}^{u})$ (7.5)
$=\delta(A_{t+1},A_{t}[1..N_{t}])\>\delta(G_{t+1},G_{t})~{}.$
In [Schatzmann:2007:NAACL], the agenda update ($A_{t}\rightarrow A_{t+1}$) is viewed as a sequence of push operations, where dialogue acts get added on top of the agenda. This is typically performed using handcrafted heuristics. For example, if a $slot_{i}=value_{i}$ provided in $a^{s}_{t}$ violates one of the goal constraints in $G_{t}$, then one of the following may be pushed onto $A_{t+1}$: NEGATE, INFORM($slot_{i}=value_{j}$), or DENY($slot_{i}=value_{i},slot_{i}=value_{j}$). The push operations are also followed by a deterministic “clean up” step, where duplicated or unnecessary dialogue acts (e.g., requests for already filled goals) are removed.
7.3.2 Data-Driven Approaches
Next, we turn to approaches to create user simulators that can be learned in a fully data-driven manner from dialogue corpora. These models can operate on the level of dialogue acts (typical of modular architectures) or directly generate natural language utterances (common in end-to-end architectures), noting that hybrid solutions also exist. It is worth pointing out that the majority of the works discussed here use simulation as an interactive environment for the training of dialogue systems, and not for evaluation, which might limit their realism [Sun:2023:TOIS].
Sequence-to-Sequence Models
Sequence-to-sequence models, originally developed for machine translation, have been successfully applied to simulate user behaviour in TOD. This is a natural fit, as both tasks involve generating a sequence of tokens in response to an input sequence. Here, the input sequence is some representation of the dialogue history, and the output sequence is the next user action and/or utterance; see Table 7.1 for an overview.
Table 7.1: Overview of user simulators using a sequence-to-sequence model structure. The output may be a dialogue act (act), utterance (utt), user satisfaction score (sat), or some combination of these.
Reference Architecture Input Output Modeling Multi-
goal? domain?
[ElAsri:2016:Interspeech] RNN-LSTM feature vect. act Y N
[Crook:2017:Interspeech] RNN-GRU/LSTM utt utt N N
[Kreyssig:2018:SIGDIAL] RNN-LSTM feature vect. utt Y N
[Gur:2018:SLT] RNN-GRU act act Y N
[Lin:2021:SIGDIAL] Transformer feature vect. act Y Y
[Sun:2021:SIGIR] Transformer context act+sat Y Y
[Lin:2022:SIGDIAL] Transformer context act+utt Y Y
[Kim:2022:SIGIR] Transformer context act+utt+sat Y Y
[Sun:2023:TOIS] Transformer context act+sat Y Y
ElAsri:2016:Interspeech propose an LSTM-based approach that takes as input a sequence of dialogue contexts and outputs a sequence of user actions. At the beginning of each dialogue, a user goal with constraints and requests, $G=(C,R)$, is randomly generated, similar to the agenda-based approach, which remain fixed during the dialogue. In each turn $t$, the simulated user takes as input the conversation contexts observed so far, $\langle c_{1},\dots,c_{t}\rangle$. Specifically, $c_{t}$ consists of four components (all represented as binary vectors): the most recent machine action, the inconsistency between machine information and user goal (i.e., slots that have been misunderstood by the system so that these may be corrected), the constraint status (to inform the system about preferences), and the request status (to keep track of requests that have not yet been fulfilled). This is passed to an RNN, which encodes this dialogue history into a single vector $v_{t}$. Then, $v_{t}$ is used to initialize the decoder RNN at each time step, which outputs a sequence of $l$ user intents $\langle I_{1},\dots,I_{l}\rangle$. These intents are turned to dialogue acts $a_{t+1}^{u}$ by looking at the current user goal and uniformly drawing among the constraints/requests left. Note that the model here can be considered partially interpretable, as the factors influencing the user’s next action are made explicit. The actual decision-making, however, is governed by a machine-learned model.
Unlike the above model that works on the semantic level of dialogue acts, sequence-to-sequence models have also been used to generate the text of user utterances directly. Crook:2017:Interspeech extend a basic encoder-decoder neural network with an additional dense layer between the encoder and decoder for better representation of context, following a similar architecture that has been proposed for non-goal-driven dialogue systems in [Serban:2016:AAAI].
Kreyssig:2018:SIGDIAL present the Neural User Simulator (NUS), which follows a very similar architecture and context representation as ElAsri:2016:Interspeech, but the output is a sequence of words. NUS allows for goal changes (when the system indicates that the goal is unachievable) and is able to generate a diverse set of utterances in the same dialogue context.
Gur:2018:SLT propose an end-to-end hierarchical user simulator (HUS), which encodes the entire dialogue history based on the user goal and system dialogue act, without requiring feature extraction or a supervised signal for state tracking. A pivotal aspect of their approach is to represent system dialogue acts on a more coarse level, by replacing specific slot values with one of the following: Requested, if the value is requested by the system, ValueInGoal, if the value appears in the user goal, ValueContradictsGoal, if the value contradicts the user goal, DontCare, if the value in the user goal is flexible, or Other otherwise. The actual values can then be substituted either from the user goal, system dialogue act, or from the knowledge base. A modified version of their model, the variational hierarchical user simulator (VHUS), introduces a variational sampling step before decoding the user turn to generate more diverse responses. Gur:2018:SLT also develop a goal regularization approach to enforce a more direct correspondence and avoid generating too long dialogues when the user diverges from the initial goal.
Transformer-based models [Vaswani:2017:NIPS], renowned for their strong natural language understanding capabilities, have spurred a new generation of user simulation approaches. Lin:2021:SIGDIAL present a transformer-based user simulator (TUS), employing a feature-based input representation similar to ElAsri:2016:Interspeech, but design it to be domain-independent. They use variable input lengths based on the slots mentioned by the system. Using a transformer model as the user policy network, input sequences of arbitrary lengths are handled and the relationships between the different slots can be captured by self-attention. The output includes the domain that should be selected for the given turn and a list of one-hot vectors to determine the values of the slots, using a coarse-grained representation similar to Gur:2018:SLT. (Note that the user intent is not explicitly predicted, but since only two possible user intents are considered, INFORM and REQUEST, those can be inferred based on how slot-values are represented.)
Sun:2021:SIGIR propose to simulate not only the next user action but also predict user satisfaction. The dialogue context, simply a concatenation of all utterances, is encoded into a latent representation using BERT [Devlin:2019:NAACL]. Two independent models are trained for two tasks, such that “action prediction takes the predicted output of satisfaction prediction model as the input” [Sun:2021:SIGIR].
Lin:2022:SIGDIAL present a generative transformer-based simulator (GenTUS) that generates semantic as well as natural language representations of the user action, that is, both a dialogue act and an utterance. Specifically, GenTUS takes the last system dialogue act, the previous three user dialogue acts, the user goal, and the turn number $t$ as input (encoded as a JSON-formatted string). The decoder generates the next user dialogue act and the corresponding language response as output, leveraging the ontology to constrain the generation process to prevent generating illegal semantic actions.
Kim:2022:SIGIR fine-tune a T5 model [Raffel:2020:JMLR] in a multi-task learning setting to create a user simulator that predicts user satisfaction and actions, and also generates utterances. Through an ablation study, the authors show that the three tasks help each other to better simulate users.
Sun:2023:TOIS propose a metaphorical user simulator, MetaSim, which aims to improve the realism of simulation by mimicking the user’s analogical thinking in interactions with the system. It relies on a database of training dialogue records, from which it retrieves the top-$k$ most similar utterances based on the current dialogue state, to serve as the metaphor. The metaphor is then taken as additional input to the policy, to determine the next user action, along with a prediction of user satisfaction. Both the metaphor module and the dialogue policy are based on T5.
Summary. Looking at Table 7.1, we can observe a trend of moving from manual feature engineering to progressively adopting end-to-end approaches. Further, approaches have become not only multi-domain, but also multi-task, simultaneously training models to perform multiple tasks relevant to user simulation, such as dialogue act prediction and utterance generation or user satisfaction estimation. It is interesting to observe that while early work was done exclusively in the field of dialogue systems [Schatzmann:2006:KER], more recent work is being increasingly carried out within the field of natural language processing and, to a smaller extent, within information retrieval. As models become more end-to-end, their interpretability diminishes, while offering limited control over the behaviour of the simulated user. Effectively, dialog planning can only be controlled indirectly, through the input data that is supplied for training, which is problematic for several reasons. First, it requires a sufficient amount of training data in the target domains, which makes these approaches less effective in low-resource scenarios. Second, if some actions or conversation paths are not present in the training corpus then those will not be captured in the simulator. Third, data-driven simulators may suffer from bias present in the corpus. Fourth, if users with various personalities are to be modeled (e.g., based on the level of cooperativeness), these characteristics need to be provided as supervised training signals along with each training example [Gur:2018:SLT]. Different experiments might require additional data annotation/collection effort, which would contradict the primary goal of using simulation as a means of automatic evaluation without human involvement. Generally, basing user dialogue policy on sequence-to-sequence models suffers from the same fundamental issues as end-to-end simulators; cf. our earlier discussion in Section 7.2.2.
In-Context Learning
Recently, the emergence of large pre-trained language models capable of following instructions have been utilized for creating user simulators through a technique known as in-context learning. This approach involves providing the LLM with a prompt that describes the task, optionally accompanied by a few examples for demonstrating the desired behaviour, to guide its responses, thus eliminating the need for fine-tuning on historical dialogues. This is a rapidly evolving area of research, with new approaches and techniques continuing to emerge. Rather than attempting an exhaustive review, we highlight some key research themes and discuss some of the open challenges.
In-context learning has been shown to be effective in simulating user utterances [Davidson:2023:arXiv, Terragni:2023:arXiv] and feedback [Hu:2023:CIKM]. It has also been successfully applied to simulate both the user and system sides of a conversation, generating annotated dialogues [Li:2022:EMNLP]. Within this line of research, several avenues have been explored, including investigating strategies for sampling dialogue examples as “shots” for the model, with techniques ranging from random sampling [Li:2022:EMNLP] to similarity-based selection according to the user goal [Terragni:2023:arXiv]. Additionally, researchers have experimented with various prompt design strategies, such as varying the number of examples and comparing structured (bullet point) formats with descriptive ones [Terragni:2023:arXiv], as well as utilizing different LLMs for the task [Terragni:2023:arXiv, Davidson:2023:arXiv].
These approaches are able to generate linguistically diverse and human-like utterances, but tend to be sensitive to the prompt used [Davidson:2023:arXiv]. Compared to sequence-to-sequence approaches, there is even less control over the data generation process as there are no guarantees that the LLM would follow the instructions in the prompt. For example, Terragni:2023:arXiv find that “in contrast to an agenda-based simulator, a prompt-based user simulator may easily give up after a dialog breakdown, deviating from the requirement to fulfill all the objectives described in the prompt.” It is also common for LLMs to hallucinate, e.g., “the user arbitrarily expresses disinterest in Chinese food, despite it not being mentioned in their original goal” [Terragni:2023:arXiv]. Additionally, these types of simulators are prone to getting stuck in loops with the TOD system, wherein both keep repeating their last utterance [Davidson:2023:arXiv, Terragni:2023:arXiv]. Terragni:2023:arXiv further observe cases where the user simulator disrupts the dialog by assuming the role of an assistant and conclude that “the issue likely arises from the training data for LLMs, which typically emphasizes emulating assistant behaviour rather than accurately representing user behaviour.” Mitigating these deficiencies of LLMs is an active area of ongoing work. For example, Li:2022:EMNLP propose a method to generate a belief state at each turn, to check whether it is consistent with the user utterance. Further improvements proposed to prompt design include “prompting the user simulator to repeat the sub-goal on which it is grounding its output” to provide additional grounding and “recasting the dialogue as a chain-of-thought reasoning problem” [Davidson:2023:arXiv].
7.4 From Task-Oriented Dialogue to Conversational Information Access
Next, we turn our attention to conversational information access, which represents a particular kind of task-oriented dialogue. It is a broad task that encompasses the goals of conversational search, recommendation, and question answering. CIA systems are expected to (1) support multiple user goals, including search, recommendation and exploratory information gathering, (2) learn user preferences and personalize responses accordingly, and (3) be capable of taking initiative [Balog:2021:DESIRES]. Radlinski:2017:CHIIR identify desirable properties of conversational search system, which include the following: (1) user revealment (enabling the user to express/discover their information need), (2) system revealment (enabling the system to reveal its capabilities and corpus), (3) mixed initiative (allowing both the user and system to take initiative), (4) memory (the ability to reference and incorporate past statements), and (5) set retrieval (considering utility over sets of complementary items). It is important to note that while the goals that CIA systems should support are clear, much of the envisioned functionality, as of today, mostly exists only on a conceptual level and approaches that support multiple conversational goals (search, recommendation, and QA) are yet to be developed. It is only very recently that datasets that support multiple conversational goals have started to become available [Bernard:2023:SIGIR]. In existing work, simulation approaches are developed in a goal-specific manner, therefore we will discuss them along this distinction: conversational search (Section 7.5) and conversational recommendation (Section 7.6). [^7] Conversational QA is not covered as there has not been simulation work on that until very recently [Abbasiantaeb:2024:WSDM].
Nevertheless, it is helpful to conceptualize the process of information access as a type of task-oriented dialogue, and abstract out both low-level actions and decision points (Section 7.4.1) as well as higher-level dialogue patterns (Section 7.4.1) that manifest during conversational interactions.
7.4.1 Action Space